# Sketch-guided Cage-based 3D Gaussian Splatting Deformation

### v2

@@ Abstract
3D Gaussian Splatting (GS) is one of the most promising novel 3D representations that has received great interest in computer graphics and computer vision. While various systems have introduced editing capabilities for 3D GS, such as those guided by text prompts, fine-grained control over deformation remains an open challenge. In this work, we present a novel sketch-guided 3D GS deformation system that allows users to intuitively modify the geometry of a 3D GS model by drawing a silhouette sketch from a single viewpoint. Our approach introduces a new deformation method that combines cage-based deformations with a variant of Neural Jacobian Fields, enabling precise, fine-grained control. Additionally, it leverages large-scale 2D diffusion priors and ControlNet to ensure the generated deformations are semantically plausible. Through a series of experiments, we demonstrate the effectiveness of our method and showcase its ability to animate static 3D GS models as one of its key applications.
Figure 1: Deformations of 3D Gaussian splats (GS) using our sketch-guided deformation method. Given a 3D GS scene, the user can deform the 3D GS by drawing a deformed silhouette sketch of a single view. (A) and (B) show examples with synthetic data and real-world large-scale data respectively. We can also produce an animation of a of 3D GS by few ($\geq 2$) keyframe sketches, as shown in (C).
@@ 1 Introduction
Figure 2: Overview of our method. We start with a 3D GS model (A). Users can select one specific sketch view (B) to draw an edited silhouette sketch (C). This user-drawn sketch (C) and rendering from the chosen view (B) are fed into ControlNet [39] resulting in a deformed reference image (D). We then optimize the cage of the source 3D GS model (A) using two losses: (1) a silhouette loss $L_{sil}$ between the silhouette (E) of the deformed GS model (F) and the silhouette of the generated reference image (D), and (2) a 3D-aware SDS loss from $4$ random views (G) conditioned on the reference image (D).
Editing of 3D models and shapes often arises in computer graphics, computer animation, and geometric modeling. This editing is usually carried out by deforming the 3D models. In this work, we aim to provide such editing-through-deformation capabilities for one of the most promising novel representations of 3D geometry - 3D Gaussian Splatting (GS) [16], which offers great real-time novel view rendering ability and better photorealistic reconstruction than previous methods.
Similarly to other geometric representations such as NeRFs [24] and triangle meshes, various control mechanisms were proposed for editing GS, such as text prompts [5, 34, 4, 35] and video priors [28, 20]. Unfortunately, these types of controls are designed for broad, high-level edits (ones within the capabilities of a novice user), without enabling fine-grained control over the deformation. On the other hand, some investigation has been made into more direct, geometrical editing, e.g., via physics-based simulation [37] - this again offers limited editing capabilities.
@@
To tackle these issues, in this work, we introduce the first sketch-guided 3D GS deformation system, which enables the user to intuitively interact with a simple 2D sketch of the object, and by which induces a 3D deformation of the Gaussians. To achieve this, we propose several technical contributions: 1) geometrically, to ensure the deformations produced are regulated, we propose a novel deformation framework for GS, based on cage-based deformations, which are in turn controlled by deformation Jacobians [1]. 2) Semantically, we leverage ControlNet [39] and Score distillation Sampling(SDS) [27] to ensure a semantically-meaningful, plausible 3D GS deformation. Together, these two contributions enable the user to deform the shape freely, while preserving its integrity, see Figure 5.
@@ 2.1 Sketch-based 3D shape Editing
Sketching is a widely used modeling paradigm in geometric modeling and computer graphics. Early methods such as Teddy [13] and Fibremesh [26] construct smooth shapes guided by user-specified 2D sketches. After generating the initial shape, deformations can be done by drawing reference strokes or deforming the generated curve.
Some works [25, 17, 42, 19] focus on the sketch-based deformation that uses individual drawn strokes as a deformation clue. (For a thorough review of the 3D shape modeling, we refer to [6].) All these methods combine the user constraints from sketches with some geometric regularizer such as the Laplacian. Whereas, these geometric energies were designed to preserve certain properties, such as smoothness, and couldn’t take into account the semantic information of the object. This can sometimes lead to unnatural deformation, e.g. bending the straight shape.
More recently, data-driven methods were applied to sketch-based 3D shape modeling [9, 23, 21, 3]. Some works [9, 3] trained neural networks to generate 3D meshes conditioned on the input sketch. These methods always need large-scale datasets to train the networks and the editing can only be done for generated shapes. For more general sketch-based editing,  [23, 21] edited the shapes (represented by Neural Radiance Field) by 2D sketch matching and employed Score Distillation Sampling(SDS) [27] to produce natural-looking editing that satisfied the semantic of the text prompt.
@@ 2.2 SDS in 3D shape Editing
Score Distillation Sampling(SDS) was first introduced in  [27] as a 3D shape generation method based on 2D diffusion prior. The core idea of SDS is to make the renderings of generated 3D shapes look natural from any random viewpoint. Since it is an image-based score, it can be applied to 3D editing of any representation, e.g. triangular mesh [36, 38], NeRF [23] and 3D GS [20]. However, the original SDS used a 2D image diffusion model with only 2D knowledge, leading to the view inconsistency problem of SDS. Recently, by using 3D data in training the diffusion model, Multi-view diffusion [29, 22, 12] was introduced to generate 3D shapes with better geometric consistency. By replacing the image diffusion model in the SDS with multi-view Diffusion, the cross-view consistency can be improved in the 3D shape editing [28, 21].
@@ 2.3 Editing 3D Gaussian Splatting
Neural Radiance Field(NeRF) [24] and 3D Gaussian Splating(3D GS) [16] are new 3D representations designed for novel-view synthesis that the 3D scene can be reconstructed from a set of images. Some work has been done for NeRF editing [23, 10, 30, 41]. Since our work focuses on sketch-based deformation of 3D GS, we will talk in more detail about editing 3D GS. [34] introduce a method that edits the 3D Gaussian Splatting(GS) scene with text instruction, powered by LLM and 2D image diffusion prior, which can achieve object texture editing and environment changing. [5] also uses 2D image diffusion prior as the guidance of the editing but introduced the Hierarchical GS to improve the editing quality. It enables object removal and addition by employing inpainting techniques.
@@
Instead of guiding the editing by 2D image diffusion prior, [4, 35] introduced methods that edit the rendered image of original 3D GS from multi-views with consistency control, and fit the changes in the edited images to 3D GS directly, which improve the efficiency and quality of editing significantly.
@@
Some works focus on the deformation of the 3D GS. Align-Your-Gaussians(AYG) [20] and DreamGaussian4D [28] introduced methods that animate a static 3D GS object to a 4D GS sequence. AYG [20] employs Video Diffusion prior to drive the temporal deformation between frames and using 2D Diffusion prior for every frame respectively to constrain the deformation validly. Instead of using Video Diffusion prior, DreamGaussian4D [28] uses a reference video in one specific view to animate the static 3D GS and applies 3D-aware Score Distillation Sampling(SDS) to propagate the deformation in every frame. The reference video was generated from an image-to-video Diffusion model based on the rendered image of static 3D GS from that specific view. Compared to AYG, DreamGaussian4D is more efficient but limited by the generation quality and universality.
@@
PhysGaussians [37] seamlessly integrates physically grounded Newtonian dynamics within 3D GS to achieve high-quality novel motion synthesis. It adapts the Material Point Method(MPM) to 3D GS which enriches 3D GS with meaningful kinematic deformation and mechanical stress attributes.
@@
SuGaR [8] employs new training energies based on original 3D GS [16] that produce 3D GS with better surface alignment and more even density distribution. These new properties make it possible to extract mesh from 3D GS only using traditional methods, such as Poisson reconstruction [15]. Given the extracted mesh, The 3D GS can be bound to the mesh and deformed by mesh deformation algorithms, such as ARAP [31]. Similar to SuGaR, [7] also proposes to bind the 3D GS kernels to the mesh, which is reconstructed by the existing methods from the input multi-view images directly. However, the result significantly depends on the extracted mesh, which can fail in scenes with complex geometry and transparent components.
@@ 3 Method
We next detail the various components of our framework, starting with an overview of the representation of 3D Gaussians [16], moving on to our cage-based deformation through jacobians, and concluding with applying this deformation technique using sketches and Score Distillation Sampling [27].
@@ 3.2 A Regularized Framework for Deforming 3D Gaussians
Deforming 3D Gaussians entails assigning a new centroid position $\mu$ and a new covariance matrix $\Sigma$ to each Gaussian. However, we wish to regulate the space of possible deformations, to avoid Gaussians floating apart, and exposing only meaningful deformations of the object the Gaussians represent. Towards this goal, we design a novel, tailor-made deformation scheme, incorporating two components: 1) a cage-based deformation [2], tailored to Gaussian Splats; 2) a method to control this cage deformation, inspired by neural jacobian fields [1]. We detail these two components next.
@@ 3.2.1 Cage-Based Deformation of Gaussians
A cage-based deformation uses a triangular mesh $S$ with vertices $V$ and triangles $T$. The mesh is deformed into another state, $S^{\prime}$, by moving its vertices. By that, the mesh defines a deformation $f_{S^{\prime}}:\mathbb{R}^{3}\to\mathbb{R}^{3}$, mapping every point in $\mathbb{R}^{3}$ as a function of the positioning of the vertices of $S^{\prime}$. This enables exposing a more meaningful, low-dimensional deformation space, controlled by the cage’s vertices. There are many different approaches to define this function; we choose to use [2] - see the supplementary material for the full details.
@@ 3.2.2 Controlling Cages via Neural Jacobian Fields
Figure 5: Comparison of using Cage or MLP+HexPlane [28] to represent deformation of the $3D$ Gaussians. MLP+Hexplane can cause severe fuzzy rendering of the deformed 3D GS because of the lack of geometry regularization. Our cage-based method produces almost lossless deformation regarding the rendering visual quality.
While the cage-based deformation already regularizes the deformation of the Gaussians, We found that optimizing directly on the cage vertices tends to produce entanglement and unsmooth cage which can lead to artifacts in the rendering of $3D$ GS, as in figure  8. We thus control the cage vertices’ position using Neural Jacobian fields (NJF) [1]. In short, NJF positions a mesh’s vertices $V^{\prime}$ from given per-triangle matrices $M_{i}\in\Re^{3\times 3}$, by minimizing the squared error between those matrices and the mesh’s per-face Jacobians $J_{i}\in\Re^{3\times 3}$, defined as
@@
We represent the rotational component using the smooth 6-DoF representation illustrated in [40] and the stretching component as a symmetric 3 by 3 matrix with 6 degrees of freedom. Since we actually have more freedom after decomposing the Jacobian field for optimization, we found that it can produce lower energy values than optimizing the Jacobian field directly.
@@
We use Score Distillation Sampling(SDS) [27] to guide our deformation to be plausible. In short, SDS renders a 3D model from different view points, and then leverages a trained 2D image diffusion model, and backpropagates the diffusion process through the differentiable renderer, to the degrees of freedom of the 3D model. we further make use of a 3D-aware image diffusion model [22] to enable a more accurate 3D consistency of the generated images. In short, this model can be conditioned on a specific viewing direction, and produces more consistent images of the same object from different view point. See Supplemental for the full details.
@@ 3.4 Sketch-guided 3D GS Deformer
The overview of the editing pipeline is shown in Figure 2. The user first selects a viewpoint $vp$ (blue camera). This viewpoint is used to extract a silhouette of the object. We render the $3D$ GS from viewpoint $vp$ to get an image $I_{vp}$ (Figure 2 B). The user additionally deforms the silhouette into $S_{vp}$ (Figure 2 C). To obtain the deformed image $I^{CN}_{vp}$ (Figure 2 D) guided by the user’s sketch, the rendering $I_{vp}$ is fed into an image-to-image diffusion model conditioned on the $S_{vp}$ by using ControlNet [39]. Our loss measures the silhouette difference $\mathcal{L}_{sil}$ between $I^{CN}_{vp}$ and $I^{def}_{vp}$ (Figure 2 E), the rendering of deformed 3D GS from viewpoint $vp$:

### v3

@@ Abstract
3D Gaussian Splatting (GS) is one of the most promising novel 3D representations that has received great interest in computer graphics and computer vision. While various systems have introduced editing capabilities for 3D GS, such as those guided by text prompts, fine-grained control over deformation remains an open challenge. In this work, we present a novel sketch-guided 3D GS deformation system that allows users to intuitively modify the geometry of a 3D GS model by drawing a silhouette sketch from a single viewpoint. Our approach introduces a new deformation method that combines cage-based deformations with a variant of Neural Jacobian Fields, enabling fine-grained control. Additionally, it leverages 2D diffusion priors and ControlNet to ensure the generated deformations are semantically plausible. Through a series of experiments, we demonstrate the effectiveness of our method and showcase its ability to animate static 3D GS models as one of its applications.
Figure 1: Deformations of 3D Gaussian splats (GS) using our sketch-guided deformation method. Given a 3D GS scene, the user can deform the 3D GS by drawing a deformed silhouette sketch of a single view. (A) and (B) show examples with synthetic data and real-world large-scale data, respectively. We can also produce an animation of a of 3D GS by few ($\geq 2$) keyframe sketches, as shown in (C).
@@ 1 Introduction
Figure 2: Overview of our method. We start with a 3D GS model (A). Users can select one specific sketch view (B) to draw an edited silhouette sketch (C). This user-drawn sketch (C) and rendering from the chosen view (B) are fed into ControlNet [zhang2023adding] resulting in a deformed reference image (D). We then optimize the cage of the source 3D GS model (A) using two losses: (1) a silhouette loss $L_{sil}$ between the silhouette (E) of the deformed GS model (F) and the silhouette of the generated reference image (D), and (2) a 3D-aware SDS loss from $4$ random views (G) conditioned on the reference image (D).
Editing of 3D models and shapes often arises in computer graphics, computer animation, and geometric modeling. This editing is usually carried out by deforming the 3D models. In this work, we aim to provide such editing-through-deformation capabilities for one of the most promising novel representations of 3D geometry - 3D Gaussian Splatting (GS) [kerbl20233d], which offers great real-time novel view rendering ability and better photorealistic reconstruction than previous methods.
Similarly to other geometric representations such as NeRFs [mildenhall2020nerf] and triangle meshes, various control mechanisms were proposed for editing GS, such as text prompts [chen2024gaussianeditor, wang2024gaussianeditor, chen2024dge, wu2024gaussctrl] and video priors [ren2023dreamgaussian4d, ling2024align]. Unfortunately, these types of controls are designed for broad, high-level edits (ones within the capabilities of a novice user), without enabling fine-grained control over the deformation. On the other hand, some investigation has been made into more direct, geometrical editing, e.g., via physics-based simulation [xie2024physgaussian] - this again offers limited editing capabilities.
@@
To tackle these issues, in this work, we introduce the first sketch-guided 3D GS deformation system, which enables the user to intuitively interact with a simple 2D sketch of the object, and by which induces a 3D deformation of the Gaussians. To achieve this, we propose several technical contributions: 1) geometrically, to ensure the deformations produced are regulated, we propose a novel deformation framework for GS, based on cage-based deformations, which are in turn controlled by deformation Jacobians [aigerman2022neural]. 2) Semantically, we leverage ControlNet [zhang2023adding] and Score distillation Sampling(SDS) [poole2022dreamfusion] to ensure a semantically-meaningful, plausible 3D GS deformation. Together, these two contributions enable the user to deform the shape freely, while preserving its integrity, see Figure 5.
@@ 2.1 Sketch-based 3D shape Editing
Sketching is a widely used modeling paradigm in geometric modeling and computer graphics. Early methods such as Teddy [igarashi2006teddy] and Fibremesh [nealen2007fibermesh] construct smooth shapes guided by user-specified 2D sketches. After generating the initial shape, deformations can be done by drawing reference strokes or deforming the generated curve.
Some works [nealen2005sketch, kho2005sketching, zimmermann2008sketching, Kraevoy:09] focus on the sketch-based deformation that uses individual drawn strokes as a deformation clue. (For a thorough review of the 3D shape modeling, we refer to [ding2016survey].) All these methods combine the user constraints from sketches with some geometric regularizer such as the Laplacian. Whereas, these geometric energies were designed to preserve certain properties, such as smoothness, and couldn’t take into account the semantic information of the object. This can sometimes lead to unnatural deformation, e.g. bending the straight shape.
More recently, data-driven methods were applied to sketch-based 3D shape modeling [guillard2021sketch2mesh, mikaeili2023sked, liu2024sketchdream, binninger2024sens]. Some works [guillard2021sketch2mesh, binninger2024sens] trained neural networks to generate 3D meshes conditioned on the input sketch. These methods always need large-scale datasets to train the networks and the editing can only be done for generated shapes. For more general sketch-based editing,  [mikaeili2023sked, liu2024sketchdream] edited the shapes (represented by Neural Radiance Field) by 2D sketch matching and employed Score Distillation Sampling(SDS) [poole2022dreamfusion] to produce natural-looking editing that satisfied the semantic of the text prompt.
@@ 2.2 SDS in 3D shape Editing
Score Distillation Sampling(SDS) was first introduced in  [poole2022dreamfusion] as a 3D shape generation method based on 2D diffusion prior. The core idea of SDS is to make the renderings of generated 3D shapes look natural from any random viewpoint. Since it is an image-based score, it can be applied to 3D editing of any representation, e.g. triangular mesh [xie2023dragd3d, yoo2024plausible], NeRF [mikaeili2023sked] and 3D GS [ling2024align]. However, the original SDS used a 2D image diffusion model with only 2D knowledge, leading to the view inconsistency problem of SDS. Recently, by using 3D data in training the diffusion model, Multi-view diffusion [shi2023mvdream, liu2023zero, huang2024epidiff] was introduced to generate 3D shapes with better geometric consistency. By replacing the image diffusion model in the SDS with multi-view Diffusion, the cross-view consistency can be improved in the 3D shape editing [ren2023dreamgaussian4d, liu2024sketchdream].
@@ 2.3 Editing 3D Gaussian Splatting
Neural Radiance Field(NeRF) [mildenhall2020nerf] and 3D Gaussian Splating(3D GS) [kerbl20233d] are new 3D representations designed for novel-view synthesis that the 3D scene can be reconstructed from a set of images. Some work has been done for NeRF editing [mikaeili2023sked, haque2023instruct, song2023blending, zhuang2023dreameditor]. Since our work focuses on sketch-based deformation of 3D GS, we will talk in more detail about editing 3D GS. [wang2024gaussianeditor] introduce a method that edits the 3D Gaussian Splatting(GS) scene with text instruction, powered by LLM and 2D image diffusion prior, which can achieve object texture editing and environment changing. [chen2024gaussianeditor] also uses 2D image diffusion prior as the guidance of the editing but introduced the Hierarchical GS to improve the editing quality. It enables object removal and addition by employing inpainting techniques.
@@
Instead of guiding the editing by 2D image diffusion prior, [chen2024dge, wu2024gaussctrl] introduced methods that edit the rendered image of original 3D GS from multi-views with consistency control, and fit the changes in the edited images to 3D GS directly, which improve the efficiency and quality of editing significantly.
@@
Some works focus on the deformation of the 3D GS. Align-Your-Gaussians(AYG) [ling2024align] and DreamGaussian4D [ren2023dreamgaussian4d] introduced methods that animate a static 3D GS object to a 4D GS sequence. AYG [ling2024align] employs Video Diffusion prior to drive the temporal deformation between frames and using 2D Diffusion prior for every frame respectively to constrain the deformation validly. Instead of using Video Diffusion prior, DreamGaussian4D [ren2023dreamgaussian4d] uses a reference video in one specific view to animate the static 3D GS and applies 3D-aware Score Distillation Sampling(SDS) to propagate the deformation in every frame. The reference video was generated from an image-to-video Diffusion model based on the rendered image of static 3D GS from that specific view. Compared to AYG, DreamGaussian4D is more efficient but limited by the generation quality and universality.
@@
PhysGaussians [xie2024physgaussian] seamlessly integrates physically grounded Newtonian dynamics within 3D GS to achieve high-quality novel motion synthesis. It adapts the Material Point Method(MPM) to 3D GS which enriches 3D GS with meaningful kinematic deformation and mechanical stress attributes.
@@
SuGaR [guedon2024sugar] employs new training energies based on original 3D GS [kerbl20233d] that produce 3D GS with better surface alignment and more even density distribution. These new properties make it possible to extract mesh from 3D GS only using traditional methods, such as Poisson reconstruction [kazhdan2006poisson]. Given the extracted mesh, The 3D GS can be bound to the mesh and deformed by mesh deformation algorithms, such as ARAP [sorkine2007rigid]. Similar to SuGaR, [gao2024mesh] also proposes to bind the 3D GS kernels to the mesh, which is reconstructed by the existing methods from the input multi-view images directly. However, the result significantly depends on the extracted mesh, which can fail in scenes with complex geometry and transparent components.
@@ 3 Method
We next detail the various components of our framework, starting with an overview of the representation of 3D Gaussians [kerbl20233d], moving on to our cage-based deformation through jacobians, and concluding with applying this deformation technique using sketches and Score Distillation Sampling [poole2022dreamfusion].
@@ 3.2 A Regularized Framework for Deforming 3D Gaussians
Deforming 3D Gaussians entails assigning a new centroid position $\mu$ and a new covariance matrix $\Sigma$ to each Gaussian. However, we wish to regulate the space of possible deformations, to avoid Gaussians floating apart, and exposing only meaningful deformations of the object the Gaussians represent. Towards this goal, we design a novel, tailor-made deformation scheme, incorporating two components: 1) a cage-based deformation [ben2009variational], tailored to Gaussian Splats; 2) a method to control this cage deformation, inspired by neural jacobian fields [aigerman2022neural]. We detail these two components next.
@@ 3.2.1 Cage-Based Deformation of Gaussians
A cage-based deformation uses a triangular mesh $S$ with vertices $V$ and triangles $T$. The mesh is deformed into another state, $S^{\prime}$, by moving its vertices. By that, the mesh defines a deformation $f_{S^{\prime}}:\mathbb{R}^{3}\to\mathbb{R}^{3}$, mapping every point in $\mathbb{R}^{3}$ as a function of the positioning of the vertices of $S^{\prime}$. This enables exposing a more meaningful, low-dimensional deformation space, controlled by the cage’s vertices. There are many different approaches to define this function; we choose to use [ben2009variational] - see the supplementary material for the full details.
@@ 3.2.2 Controlling Cages via Neural Jacobian Fields
Figure 5: Comparison of using Cage or MLP+HexPlane [ren2023dreamgaussian4d] to represent deformation of the $3D$ Gaussians. MLP+Hexplane can cause severe fuzzy rendering of the deformed 3D GS because of the lack of geometry regularization. Our cage-based method produces almost lossless deformation regarding the rendering visual quality.
While the cage-based deformation already regularizes the deformation of the Gaussians, We found that optimizing directly on the cage vertices tends to produce entanglement and unsmooth cage which can lead to artifacts in the rendering of $3D$ GS, as in figure  8. We thus control the cage vertices’ position using Neural Jacobian fields (NJF) [aigerman2022neural]. In short, NJF positions a mesh’s vertices $V^{\prime}$ from given per-triangle matrices $M_{i}\in\Re^{3\times 3}$, by minimizing the squared error between those matrices and the mesh’s per-face Jacobians $J_{i}\in\Re^{3\times 3}$, defined as
@@
We represent the rotational component using the smooth 6-DoF representation illustrated in [zhou2019continuity] and the stretching component as a symmetric 3 by 3 matrix with 6 degrees of freedom. Since we actually have more freedom after decomposing the Jacobian field for optimization, we found that it can produce lower energy values than optimizing the Jacobian field directly.
@@
We use Score Distillation Sampling(SDS) [poole2022dreamfusion] to guide our deformation to be plausible. In short, SDS renders a 3D model from different view points, and then leverages a trained 2D image diffusion model, and backpropagates the diffusion process through the differentiable renderer, to the degrees of freedom of the 3D model. we further make use of a 3D-aware image diffusion model [liu2023zero] to enable a more accurate 3D consistency of the generated images. In short, this model can be conditioned on a specific viewing direction, and produces more consistent images of the same object from different view point. See Supplemental for the full details.
@@ 3.4 Sketch-guided 3D GS Deformer
The overview of the editing pipeline is shown in Figure 2. The user first selects a viewpoint $vp$ (blue camera). This viewpoint is used to extract a silhouette of the object. We render the $3D$ GS from viewpoint $vp$ to get an image $I_{vp}$ (Figure 2 B). The user additionally deforms the silhouette into $S_{vp}$ (Figure 2 C). To obtain the deformed image $I^{CN}_{vp}$ (Figure 2 D) guided by the user’s sketch, the rendering $I_{vp}$ is fed into an image-to-image diffusion model conditioned on the $S_{vp}$ by using ControlNet [zhang2023adding]. Our loss measures the silhouette difference $\mathcal{L}_{sil}$ between $I^{CN}_{vp}$ and $I^{def}_{vp}$ (Figure 2 E), the rendering of deformed 3D GS from viewpoint $vp$: